Directions: Write your answer in the space provided. Show all of your work in order to receive credit for your answer for tho Is this correct?
Answer #1

Kp = kc (RT) on= change in no-onds of gas . N₂ (g) + 3 H₂(g) 2NH3(g) Ke=4.58104 T= 4ook R=0.08206 Lam molk Dn=2-(3+1) Kp = Kc